OpenStudy (anonymous):

Find the measure of the three angles of a triangle if the second angle is 8 degrees less than three times the first and the third is one degree more than seven times the first .

OpenStudy (gorv):

6/16

OpenStudy (gorv):

HHTT+HTHT+HTTH+THHT+THTH+TTHH

OpenStudy (gorv):

@janhell123 GOT IT???

OpenStudy (gorv):

sry

OpenStudy (anonymous):

i don't get it

OpenStudy (gorv):

wait i ans other q in ur q

OpenStudy (anonymous):

post your answer here

OpenStudy (anonymous):

and solution

OpenStudy (gorv):

angles a1 a2 a3 a2=3a1-8 a3=7a1+1 in triangle a1+a2+a3=180 put values of a2 and a3 a1+3a1-8+7a1+1=180 11a1=187 a1=17 a2=3*17-8=43 a3=17*7+1=120

OpenStudy (gorv):

@janhell123

OpenStudy (gorv):

give medal if u got it